OpenStudy (anonymous):

when you estiamte 1 does it can 0 ever be an answer

OpenStudy (anonymous):

like what?

OpenStudy (anonymous):

can a 0 be an answer when you estimate 1

OpenStudy (anonymous):

like .1 or 1.0 ?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

1

OpenStudy (anonymous):

i do not under stand

OpenStudy (qwertty123):

@mathstudent55

OpenStudy (anonymous):

i do not under stand

OpenStudy (anonymous):

t) Which statement about the best estimation strategy for this problem is correct? Ms. Levy paid $189 for a hotel room and she bought two concert tickets priced at $99 each. How much money did Ms. Levy spend altogether? A student answered that Ms. Levy spent $576. A. ($200 + $100) × 2 = t The student's answer was reasonable. B. $200 + (2 × $100) = t The student's answer was not reasonable. C. $200 + ($100 ÷ 2) = t The student's answer was not reasonable. D. ($200 × 2) + $100 = t The student's answer was not reasonable. t) Which statement about the best estimation strategy for this problem is correct? Ms. Levy paid $189 for a hotel room and she bought two concert tickets priced at $99 each. How much money did Ms. Levy spend altogether? A student answered that Ms. Levy spent $576. A. ($200 + $100) × 2 = t The student's answer was reasonable. B. $200 + (2 × $100) = t The student's answer was not reasonable. C. $200 + ($100 ÷ 2) = t The student's answer was not reasonable. D. ($200 × 2) + $100 = t The student's answer was not reasonable. t) Which statement about the best estimation strategy for this problem is correct? Ms. Levy paid $189 for a hotel room and she bought two concert tickets priced at $99 each. How much money did Ms. Levy spend altogether? A student answered that Ms. Levy spent $576. A. ($200 + $100) × 2 = t The student's answer was reasonable. B. $200 + (2 × $100) = t The student's answer was not reasonable. C. $200 + ($100 ÷ 2) = t The student's answer was not reasonable. D. ($200 × 2) + $100 = t The student's answer was not reasonable.